# Influence of the residual magnetic field on the azimuthal distribution of final-state particles in photon-nuclear processes

### v1

@@ I Introduction
However, it is noteworthy that since electromagnetic fields act as retarded potentials, residual magnetic fields persist in the particle production region even after the collision [19, 20]. These fields deflect the trajectories of final-state particles, thereby distorting the experimentally measured azimuthal modulation patterns Consequently, this effect compromises the accuracy of extracting geometrical information of the collision system using such observables. In this paper, we calculate the coordinate-space and momentum-space distribution of photoproduced mesons in Au+Au collisions at 200 GeV and simulate their isotropic decays into pairs. Using the UrQMD model [21, 22], we simulate Au+Au collisions across different centrality classes and compute the residual magnetic fields generated by strongly interacting particles. We then apply these residual magnetic fields to the pairs and quantify the resulting cosine modulation induced by these field effects. Our calculations reveal that in peripheral collisions, the abundant particle production in the interaction region leads to a significantly slower decay of the residual magnetic field and markedly alters its spatial distribution compared to ultra-peripheral collisions. Consequently, this effect produces a pronounced impact on the azimuthal modulation pattern in such collision systems. Our study provides a new theoretical reference for future experimental investigations that aim to determine the initial geometry using polarized photons in peripheral collisions. Although this paper primarily investigates the influence of the residual magnetic field on the angular distribution of final-state particles in photon-nucleus interactions, such effects should also impact the angular distribution of particles produced via photon-photon interactions[23, 24]. Therefore, corresponding contributions must be considered in related studies.
@@ II Methodology
The computational approach adopted in this paper is as follows: First, we calculate the coordinate-space and momentum-space distribution of photoproduced mesons. The mesons are sampled from these distributions and decay isotropically (without considering polarization effects, focusing solely on the potential influence of the residual magnetic field), with their finite lifetime properly considered. Next, we use the UrQMD model and Electrodynamics to compute the spatial and temporal evolution of the magnetic field for different collision centralities. Finally, we apply the magnetic field to the pairs from decay and calculate the second-order cosine modulation coefficient of their final-state azimuthal distribution. In this study, the right-handed coordinate system for the vector meson decay is constructed such that the z-axis points along the vector meson’s flight direction in the photon-nucleon center-of-mass frame, the y-axis is normal to the photoproduction plane, and the x-axis is given by y × z to complete the right-handed system, is the azimuthal angle.
@@ II.2 The calculation of residual magnetic fields
The total magnetic field is the sum of the magnetic filed induced by the two colliding Au nuclei.
@@ III Results
We investigate the azimuthal modulation (2) of final-state particles by applying the magnetic field to the decays from the photoproduced mesons which are sampled according to the calculated distribution described in Section 2.1, where is defined in rest frame as . We selected three characteristic time points to monitor the evolution of 2: (1) fm/, corresponding to the instant of nuclear collision with initial photoproduction. (2) fm/,when most mesons have decayed and no additional particles are produced in the nuclear overlap region (particularly for peripheral collisions). (3) fm/, representing the final-state configuration where the decayed pairs have propagated sufficiently through the magnetic field. Figure 4 shows the 2 induced by the residual magnetic field in UPC case as a function of for different impact parameter ranges. It can be seen that for the UPC case, 2 does not show significant changes in different b ranges with and time. This is because, as shown by the blue curve in Fig. 3, the magnetic field generated after the ”collisions” of the two Au nuclei decays rapidly and is primarily concentrated near the origin, where very few are produced. Consequently, it does not exhibit a significant deflection effect on the final-state particles. Therefore, in UPC case, the 2 modulation can be solely attributed to the linear polarization of photons.

### v2

@@ I Introduction
However, it is noteworthy that since electromagnetic fields act as retarded potentials, residual magnetic fields persist in the particle production region even after the collision [19, 20]. These fields deflect the trajectories of final-state particles, thereby distorting the experimentally measured azimuthal modulation patterns [21, 22]. Consequently, this effect compromises the accuracy of extracting geometrical information of the collision system using such observables. In this paper, we calculate the coordinate-space and momentum-space distribution of photoproduced mesons in Au+Au collisions at 200 GeV and simulate their isotropic decays into pairs. Using the UrQMD model [23, 24], we simulate Au+Au collisions across different centrality classes and compute the residual magnetic fields generated by strongly interacting particles. We then apply these residual magnetic fields to the pairs and quantify the resulting cosine modulation induced by these field effects. Our calculations reveal that in peripheral collisions, the abundant particle production in the interaction region leads to a significantly slower decay of the residual magnetic field and markedly alters its spatial distribution compared to ultra-peripheral collisions. Consequently, this effect produces a pronounced impact on the azimuthal modulation pattern in such collision systems. Our study provides a new theoretical reference for future experimental investigations that aim to determine the initial geometry using polarized photons in peripheral collisions. Although this paper primarily investigates the influence of the residual magnetic field on the angular distribution of final-state particles in photon-nucleus interactions, such effects should also impact the angular distribution of particles produced via photon-photon interactions[25, 26]. Therefore, corresponding contributions must be considered in related studies.
@@ II Methodology
The computational approach adopted in this paper is as follows: First, we calculate the coordinate-space and momentum-space distribution of photoproduced mesons. The mesons are sampled from these distributions and decay isotropically (without considering polarization effects, focusing solely on the potential influence of the residual magnetic field), with their finite lifetime properly considered. Next, we use the UrQMD model and Electrodynamics to compute the spatial and temporal evolution of the magnetic field for different collision centralities. Finally, we apply the magnetic field to the pairs from decay and calculate the second-order cosine modulation coefficient of their final-state azimuthal distribution. In this study, the right-handed coordinate system for the vector meson decay is constructed such that the -axis points along the vector meson’s flight direction in the photon-nucleon center-of-mass frame, the -axis is normal to the photoproduction plane, and the -axis is given by to complete the right-handed system, is the azimuthal angle.
@@ II.2 The calculation of residual magnetic fields
The total magnetic field is the sum of the magnetic field induced by the two colliding Au nuclei.
@@
In addition to UrQMD, there are other models that can describe the motion of particles produced after relativistic heavy-ion collisions [43]. We compared the magnetic fileds calculated by the UrQMD model with those from the HSD model [44], shown as Fig. 4. It can be observed that at fm, fm, fm (a position located within the primary photoproduction region of , see Fig. 1), the generated from both models exhibit comparable magnitudes and temporal evolution patterns, consequently leading to similar deflection effects on the pairs originating through the photoproduction and subsequent decay of . In the present work, we will focus primarily on the results obtained from the UrQMD model calculations.
@@ III Results
We investigate the azimuthal modulation (2) of final-state particles by applying the magnetic field to the decays from the photoproduced mesons which are sampled according to the calculated distribution described in Section II. A, where is defined in rest frame as . We selected three characteristic time points to monitor the evolution of 2: (1) fm/, corresponding to the instant of nuclear collision with initial photoproduction. (2) fm/,when most mesons have decayed and no additional particles are produced in the nuclear overlap region (particularly for peripheral collisions). (3) fm/, representing the final-state configuration where the decayed pairs have propagated sufficiently through the magnetic field. Figure 5 shows the 2 induced by the residual magnetic field in UPC case as a function of for different impact parameter ranges. It can be seen that for the UPC case, 2 does not show significant changes in different b ranges with and time. This is because, as shown by the blue curve in Fig. 3, the magnetic field generated after the ”collisions” of the two Au nuclei decays rapidly and is primarily concentrated near the origin, where very few are produced. Consequently, it does not exhibit a significant deflection effect on the final-state particles. Therefore, in UPC case, the 2 modulation can be solely attributed to the linear polarization of photons.